OpenStudy (anonymous):

how to simplify x^-6-y^-6/(xy)^10

OpenStudy (anonymous):

hi are you there?

OpenStudy (anonymous):

(x^-6-y^-6)/(xy)^10 x^-6-y^-6 =---------- (xy)^10 x^-6 y^-6 =----- - ------ (xy)^10 (xy)^10 x^-6 y^-6 =-------- - ---------- note here just do addition or subtraction on like terms (x^10)y^10 (x^10)y^10 1 1 =---------- - ---------- answer...here we added (6+10) on each both on x (x^16)y^10 (x^10)y^16 and y from the denominator.